# Capturing reduced-order quantum many-body dynamics out of equilibrium via neural ordinary differential equations

### v1

@@ Abstract
Out‑of‑equilibrium quantum many‑body systems – such as multi-electron atoms and molecules driven by strong laser fields, quenched ultracold gases, and ultrafast-excited solids – exhibit rapid correlation buildup that underlies many emerging phenomena. Exact wave‑function methods to describe this scale exponentially with particle number; simpler mean‑field approaches neglect essential two‑particle correlations. The time-dependent two-particle reduced density matrix formalism offers a middle ground by propagating the two‑particle density matrix and closing the Bogoliubov–Born–Green–Kirkwood–Yvon hierarchy with a reconstruction of the three‑particle cumulant. But the validity and existence of time‑local reconstruction functionals ignoring memory effects remain unclear across different dynamical regimes. We show that a neural ordinary differential equation model trained on exact two-particle reduced density matrix data (no dimensionality reduction) can reproduce its full dynamics without any explicit three‑particle information – but only in parameter regions where the Pearson correlation between the two‑ and three‑particle cumulants is large. In contrast, in the anti‑correlated or uncorrelated regime, the neural ordinary differential equation fails, indicating that no simple time‑local functional of the instantaneous two‑particle cumulant can capture the evolution. The magnitude of the time‑averaged three‑particle‑correlation buildup appears to be the primary predictor of success: For a moderate correlation buildup, both neural ordinary differential equation predictions and existing time-dependent two-particle reduced density matrix reconstructions are accurate, whereas stronger values lead to systematic breakdowns. These findings pinpoint the need for memory‑dependent kernels in the three‑particle cumulant reconstruction for the latter regime. Our results place the neural ordinary differential equation as a model‑agnostic diagnostic tool that maps the regime of applicability of cumulant expansion methods and guides the development of non‑local closure schemes. More broadly, the ability to learn high‑dimensional reduced‑density‑matrix dynamics from limited data opens a pathway to fast, data‑driven simulation of correlated quantum matter, complementing traditional numerical and analytical techniques.
@@ 1 Introduction
The exploration of quantum many-body systems far from equilibrium has become a central theme across ultracold-atom physics, condensed-matter theory, and ultrafast spectroscopy. Rapid experimental progress now enables the precise control of isolated atomic gases [kollath2005Spin-Charge], the manipulation of correlated solids with intense light fields [buzzi2021Higgs-Mediated, giannetti2016Ultrafast], the tracking of real-time electronic motion in nanostructures [driscoll2011Laser-induced] and much more [stojchevska2014Ultrafast, caillat2005Correlated, topp2018Alloptical, vodeb2024Nonequilibrium, bedow2022Emergence, budden2021Evidence, pohl2000Towards]. These advances have revealed a broad spectrum of out-of-equilibrium phenomena (for recent reviews see e.g.  [basov_towards_2017, torre_colloquium_2021, bloch_strongly_2022]), including ultrafast switches of currents in dielectrics [schiffrin_optical-field-induced_2013, wachter2014Abinitio, schultze_controlling_2013, ossiander_speed_2022], laser-induced superconductivity [cremin_photoenhanced_2019, budden2021Evidence, buzzi2021Higgs-Mediated], and ultrafast switches to hidden quantum states [stojchevska2014Ultrafast]. Parallel developments in non-equilibrium many-body theory [cazalilla2002Time-dependent, otobe2008Firstprinciples, hochstuhl2012Timedependent, hochstuhl2014Timedependent, wachter2014Abinitio, sato2018Communication, pederson2019Symplectic, eckstein2009Thermalization, schluenzen2020achieving, joost2022Dynamically, perfetto_real-time_2022, pavlyukh_time-linear_2022, lackner2015Propagating, lackner2017Highharmonic] have highlighted the need for theoretical frameworks that faithfully model (strongly) correlated dynamics of extended systems for long periods of time while remaining computationally tractable.
@@
Accurately describing the non-equilibrium dynamics of interacting quantum many-body systems remains a central challenge because the full many-body wave function resides in a Hilbert space whose dimension grows exponentially with the particle number. Wave-function-based methods, which allow a full parametrization of the Hilbert space such as the multiconfigurational time-dependent Hartree Fock method (MCTDHF) [caillat2005Correlated, hochstuhl2014Timedependent] or matrix-product-state (MPS) techniques [Scholl2005, haegeman2016Unifying, kloss2018Timedependent, Cirac2021] provide numerically exact results but are limited to modest system sizes and short propagation times. Other approaches such as the time-dependent density-functional theory (TDDFT) [runge1984DensityFunctional, ullrich2012TimeDependent] or time-dependent Hartree–Fock (TDHF) [caillat2005Correlated] scale favorably yet are limited to weakly (TDDFT) or uncorrelated (TDHF) systems. Several approaches exist that avoid the exponential scaling of wavefunctions by using a reduced object instead. One prominent example are methods that use non-equilibrium Green’s functions (NEGF) as the central object (for a review see e.g. [stefanucci_2025]). However, NEGF methods typically feature a nonlinear scaling with time. The time-dependent two-particle reduced density matrix formalism [lackner2015Propagating, lackner2017Highharmonic, donsa2023Nonequilibrium] follows a similar spirit but features a linear scaling in time: the two-particle reduced density matrix $D_{12}(t)$ contains the exact two-particle information, and the total energy can be expressed as a functional of $D_{12}(t)$ (if the Hamiltonian only contains one- and two-body operators), thereby reducing the computational cost from exponential to polynomial scaling. Its power stems from the explicit inclusion of dynamical two-particle correlations and from a systematic closure of the Bogoliubov–Born–Green–Kirkwood–Yvon hierarchy via a reconstruction of the three-particle reduced density matrix $D_{123}(t)$ as a functional of $D_{12}(t)$. The method thus relies on approximate reconstruction functionals; present state-of-the-art schemes are time-local, which guarantees the linear scaling in time of the method, and quadratic in the two-particle cumulant $\Delta_{12}(t)$, thereby neglecting memory effects and higher-order cumulants. Consequently, the accuracy of time-dependent two-particle reduced density matrix depends on the quality of the time-local $D_{123}$ reconstruction. We would like to point out that extensions of NEGFs methods featuring a linear scaling in time [schluenzen2020achieving, joost2022Dynamically, pavlyukh_time-linear_2022] face the same problem of the necessity to reconstruct unknown components of the equations of motion in a time-local fashion. So, the deep question is, in which regimes can we find a sufficiently accurate time-local reconstruction functional? Can numerical methods shed light on this problem?
@@
Today, machine learning introduces a diverse variety of complementary approaches to classical numerical modeling. Some approaches can enable integrating domain knowledge about physical laws governing the target problems of interest. In the field of quantum physics, some examples in the existing literature model open quantum systems [Coppola2025Learning, zhang2025Neural, liu2023Predicting, Mazza2021, Chen2022, Chen2025], investigate quantum-information tasks [Chen2022, Genois2021], measure entanglement [huang2025Direct], use neural quantum states to overcome the exponential scaling in the Hilbert space [Lange2024architectures], combine machine learning with quantum computing [ye2025Simulating], and enable quantum (optimal) control [Chen2025Optimal, Norambuena2024].
@@
Dynamical processes are generally governed by differential equations and have been subject to numerous investigations using machine learning [Kidger2020a, Toth2020Hamiltonian, MorrillSKF21, Dietrich2022, Carnazza2022, Carnazza2024, Schmitt2020, Lin2022, Zhao2024, Mohseni2024, Zeng2024, An2025, Kaneko2025, Sun2025, Schmitt2025, Cemin2025]. In the present study, we focus specifically on systems whose time evolution can be described by ordinary differential equations. Neural ordinary differential equations offer a powerful venue of thinking because they combine a rudimentary simulation of propagation of time through a differential equation, with data-driven ML [chen2018Neural]. It is a promising candidate for modeling diverse time-dependent datasets in and out of science [qian2021Integrating, worsham2025guide, hoege2022Improving, hananeh2021Beyond, lee2021Parameterized, niu2024applications, sorourifar2023PhysicsEnhanced, bram2024Lowdimensional]. To our knowledge, its application in classical or quantum many-body physics is scarce. Neural ordinary differential equations have been applied to reduced representations of quantum dynamics [choi2022Learning], holographic quantum chromodynamics [Hashimoto2021], or to the Hamiltonian learning problem [Heightman2025Solving]. Other machine learning approaches have been used to investigate Markovian and non-Markovian dynamics of quantum systems [Nelson2022].
Up to now, it is an open question whether neural ordinary differential equations will be useful tools to capture complex high-dimensional dynamics of many-body systems, specifically those out of equilibrium. What are the obstacles of established many-body simulations that can be overcome by neural ordinary differential equations? Can they supersede performance? Moreover, how can one infuse prior knowledge of, e.g., symmetries and conserved quantities into specifically these machine learning models, and does this actually lead to improved model fidelity? This might not be straightforward for generic dynamical processes. Finally, can we interpret [wetzel2025interpretablemachinelearningphysics] or extract new knowledge about the target (here, the dynamical quantum system) with machine learning-based ‘simulation’?
@@
In this paper, we embark on a study of quantum many-body systems out of equilibrium, viewed through two-particle reduced density matrixs [donsa2023Nonequilibrium]. For this purpose, we investigate the dynamics of the paradigmatic Fermi-Hubbard model [Qin2021]. Depending on the parameters, it shows a broad range of dynamics, reaching from weakly to strongly correlated. We demonstrate that neural ordinary differential equations offer interesting insights: We find indications that time-local reconstructions of the three-particle reduced density matrix may only be possible in certain regimes in the parameter space. The infusion of physics knowledge is neither straightforward nor does prior physical information necessarily lead to better results.
@@
The structure of the paper is as follows: In Sec. 2 we briefly introduce the quenched Fermi-Hubbard model, which will be the system under investigation. We describe the equations of motion for the two-particle reduced density matrix and concisely revisit the cumulant reconstruction needed to close them. The neural ordinary differential equation model used in this investigation is presented together with the data preprocessing. In Sec. 3, we first identify some structure in the parameter space of the Fermi-Hubbard model. This is used to investigate the performance of the neural ordinary differential equation across the parameter plane. To improve the stability of the prediction, we enforce physical constraints in the loss. Last, we discuss the results and show perspectives for future investigations in Sec. 4.
@@ 2.1 Non-equilibrium Fermi-Hubbard model
We consider a one‑dimensional Fermi–Hubbard chain comprising $M_{s}$ physical sites together with two auxiliary sites labeled 0 and $M_{s}+1$ that are fixed to be unoccupied throughout the dynamics. Dirichlet boundary conditions are imposed at these auxiliary sites, a choice that mirrors the confinement used in quantum‑simulator experiments with trapped ultracold atoms (see, e.g., Ref. [bloch2012Quantum]). While Dirichlet boundaries are natural for the experimental setting, alternative boundary conditions (e.g., periodic) can be incorporated straightforwardly within the same formalism. The initial state is prepared in a harmonic trapping potential, which is abruptly switched off at $t=0$, after which the system evolves under the standard Fermi-Hubbard Hamiltonian [donsa2023Nonequilibrium]. The full Hamiltonian in second quantization is given by
@@
Here, $J$ denotes the hopping amplitude, $\langle i,j\rangle$ indicates nearest-neighbor pairs, $a_{i\sigma}^{(\dagger)}$ are the single-particle annihilation (creation) operators, $n_{i}^{\uparrow(\downarrow)}=a_{i\uparrow(\downarrow)}^{\dagger}a_{i\uparrow(\downarrow)}$ are the corresponding spin-resolved occupation number operators at site $i$, and $U$ characterizes the on-site interaction strength governing the correlation energy of the. The system before and after the quench can be seen in Fig. 1.
@@
We use a harmonic potential
@@
which initially traps the system in the ground state of the harmonic trap and is turned off at $t=0$. This potential quench induces the dynamics. Although this is a quite simple model, it is still complex enough to show rich dynamics.
@@ 2.2 Time-evolution of the 2RDM
The two-particle reduced density matrix is defined from the NN‑body state $|\Psi(t)\rangle$ as a trace over all but two particles
$D_{12}(t)=N(N-1)\text{Tr}_{3\ldots N}|\Psi(t)\rangle\langle\Psi(t)|\,$ (3)
with N(N1)N(N-1) being the chosen normalization. The second equation in the Bogoliubov–Born–Green–Kirkwood–Yvon hierarchy, i.e., the equation of motion for the 2RDM,
$i\partial_{t}D_{12}=[h_{1}+h_{2}+W_{12},D_{12}]+\text{Tr}[W_{13}+W_{23},D_{123}]\,$, (4)
contains the single‑particle Hamiltonian $h_{i}$, the two‑body interaction $W_{ij}$, and the three-particle reduced density matrix $D_{123}$. Closing Eq. (4) therefore requires a reconstruction functional $D^{\text{R}}_{123}[D_{12}]$. This leads to the time-dependent two-particle reduced density matrix theory introduced in Refs. [lackner2015Propagating, lackner2017Highharmonic, donsa2023Nonequilibrium].
In a total‑spin‑singlet configuration, as considered here, the evaluation of Eq. (4) requires only the mixed‑spin block $\langle j_{1}\uparrow j_{2}\downarrow|D_{12}(t)|i_{1}\uparrow i_{2}\downarrow\rangle=D^{j_{1}\uparrow j_{2}\downarrow}_{i_{1}\uparrow i_{2}\downarrow}(t)$ with $|i_{1}\uparrow i_{2}\downarrow\rangle$ two-particle product states of spin orbitals, which already encodes the complete information of the full two-particle reduced density matrix $D_{12}(t)$ [lackner2015Propagating]. For completeness we note that elements of the 2RDM can also be obtained in the second quantization formalism through $D^{j_{1}\uparrow j_{2}\downarrow}_{i_{1}\uparrow i_{2}\downarrow}(t)=\langle\Psi(t)|a_{i_{1}\uparrow}^{\dagger}a_{i_{2}\downarrow}^{\dagger}a_{j_{2}\downarrow}a_{j_{1}\uparrow}|\Psi(t)\rangle$. All other spin components can be recovered from this block by straightforward exchange or spin‑flip symmetry relations.

### v2

@@ Abstract
Out‑of‑equilibrium quantum many‑body systems – such as multi-electron atoms and molecules driven by strong laser fields, quenched ultracold gases, and ultrafast-excited solids – exhibit rapid correlation buildup that underlies many emerging phenomena. Exact wave‑function methods to describe these effects scale exponentially with particle number; simpler mean‑field approaches neglect essential particle correlations. The time-dependent two-particle reduced density matrix (TD2RDM) formalism offers a middle ground by propagating the two‑particle density matrix and closing the Bogoliubov–Born–Green–Kirkwood–Yvon (BBGKY) hierarchy with a reconstruction of the three‑particle cumulant, which carries information about three-particle correlations. But the validity and existence of time‑local reconstruction functionals ignoring memory effects remain unclear across different dynamical regimes. We show that a neural ordinary differential equation (ODE) model trained on exact two-particle reduced density matrix (2RDM) data (no dimensionality reduction) can reproduce its full dynamics without any explicit three‑particle information – but only in parameter regions where the Pearson correlation between the two‑ and three‑particle cumulants is large. In contrast, in the anti‑correlated or uncorrelated regime, the neural ODE fails, indicating that no simple time‑local functional of the instantaneous two‑particle cumulant can capture the evolution. The magnitude of the time‑averaged three‑particle‑correlation buildup appears to be the primary predictor of successful extrapolation: For a moderate correlation buildup, both neural ODE predictions and existing TD2RDM reconstructions are accurate, whereas stronger values lead to systematic breakdowns. These findings pinpoint the need for memory‑dependent kernels in the three‑particle cumulant reconstruction for the latter regime. Our results place the neural ODE as a model‑agnostic diagnostic tool that maps the regime of applicability of cumulant expansion methods and guides the development of non‑local closure schemes. More broadly, the ability to learn high‑dimensional reduced‑density‑matrix dynamics from limited data opens a pathway to fast, data‑driven simulation of correlated quantum matter, complementing traditional numerical and analytical techniques.
@@ 1 Introduction
The exploration of quantum many-body systems far from equilibrium has become a central theme across ultracold-atom physics, condensed-matter theory, and ultrafast spectroscopy. Rapid experimental progress now enables the precise control of isolated atomic gases [53], the manipulation of correlated solids with intense light fields [14, 40], the tracking of real-time electronic motion in nanostructures [35] and other manipulations of quantum systems [93, 16, 95, 99, 7, 13, 78]. These advances have revealed a broad spectrum of out-of-equilibrium phenomena (for recent reviews see e.g. [6, 32, 10]), including ultrafast switches of currents in dielectrics [83, 100, 88, 71], laser-induced superconductivity [31, 13, 14], and ultrafast switches to hidden quantum states [93]. Parallel developments in non-equilibrium many-body theory [19, 72, 45, 46, 100, 82, 75, 36, 84, 49, 76, 73, 55, 56] have highlighted the need for theoretical frameworks that faithfully model (strongly) correlated dynamics of extended systems for long periods of time while remaining computationally tractable.
@@
Accurately describing the non-equilibrium dynamics of interacting quantum many-body systems remains a central challenge because the full many-body wave function resides in a Hilbert space whose dimension grows exponentially with the particle number. Wave-function-based methods, which allow, in principle, a full parametrization of the Hilbert space such as the multiconfigurational time-dependent Hartree Fock method (MCTDHF) [16, 46] or matrix-product-state (MPS) techniques [87, 42, 52, 28] provide numerically exact results but are, in general, limited to modest system sizes and short propagation times. Other approaches such as the time-dependent density-functional theory (TDDFT) [81, 97] or time-dependent Hartree–Fock (TDHF) [16] scale favorably, yet are limited to weakly (TDDFT) or uncorrelated (TDHF) systems. Several approaches exist that avoid the exponential scaling of wavefunctions by using a reduced object instead. One prominent example are methods that use non-equilibrium Green’s functions (NEGF) as the central object (for an introduction see e.g. [92]). However, NEGF methods typically feature a nonlinear scaling with time. Another strategy is the local-information time-evolution (LITE) approach [4], which decomposes the system into subsystems at different spatial scales and evolves their reduced density matrices in parallel, systematically discarding long-range entanglement above a cutoff scale while preserving the information flow without assumptions on its microscopic details; however, its computational cost scales exponentially with the subsystem scale that controls accuracy. The TD2RDM formalism [55, 56, 34] follows a similar spirit as NEGF but features a linear scaling in time: the 2RDM $D_{12}(t)$ contains the full two-particle information, and the total energy can be expressed as a functional of $D_{12}(t)$ (if the Hamiltonian only contains one- and two-body operators), thereby reducing the computational cost from exponential to polynomial scaling. Its power stems from the explicit inclusion of dynamical two-particle correlations and from a systematic closure of the BBGKY hierarchy via a reconstruction of the three-particle reduced density matrix (3RDM) $D_{123}(t)$ as a functional of $D_{12}(t)$. The method thus relies on approximate reconstruction functionals; present state-of-the-art schemes are time-local, which guarantees the linear scaling in time of the method, and quadratic in the two-particle cumulant $\Delta_{12}(t)$, thereby neglecting memory effects and higher-order cumulants. Consequently, the accuracy of TD2RDM depends on the quality of the time-local $D_{123}$ reconstruction. We would like to point out that extensions of NEGFs methods featuring a linear scaling in time [84, 49, 73] face the same problem of the necessity to reconstruct unknown components of the equations of motion in a time-local fashion. So, the deep question is, in which regimes can we find a sufficiently accurate time-local reconstruction functional? Can numerical methods shed light on this problem?
@@
Today, machine learning (ML) introduces a diverse variety of complementary approaches to classical numerical modeling, with growing emphasis on interpretability in physics [101]. Some approaches can enable integrating domain knowledge about physical laws governing the target problems of interest. In the field of quantum physics, some examples in the existing literature model open quantum systems [30, 109, 61, 62, 23, 22], investigate quantum-information tasks [23, 39], measure entanglement [48], predict the phonon blockade effect [108], use neural quantum states to overcome the exponential scaling in the Hilbert space [58, 98], combine ML with quantum computing [105], and enable quantum (optimal) control [26, 69, 12, 107].
@@
Dynamical processes are generally governed by differential equations and have been subject to numerous investigations using ML [51, 96, 65, 33, 18, 17, 85, 60, 110, 64, 106, 3, 50, 94, 86, 20]. In the present study, we focus specifically on systems whose time evolution can be described by ODEs. Neural ODEs offer a powerful venue of thinking because they combine a rudimentary simulation of propagation of time through a differential equation, with data-driven ML [24]. It is a promising candidate for modeling diverse time-dependent datasets in and out of science [79, 104, 47, 2, 59, 68, 91, 11]. To our knowledge, its application in classical or quantum many-body physics is scarce. Neural ODEs have been applied to reduced representations of quantum dynamics [27], holographic quantum chromodynamics [43], or to the Hamiltonian learning problem [44]. Other ML approaches have been used to investigate Markovian and non-Markovian dynamics or noise of quantum systems [67, 66].
Testing for the Markov property in time series is a longstanding problem in many fields such as statistics, molecular dynamics, and quantum information theory. Classical statistical approaches based on kernel smoothing [1] or local polynomial regression of the conditional characteristic function [21] suffer from the curse of dimensionality and become unreliable in moderate- to high-dimensional settings. More recent proposals employ deep conditional generative models such as mixture density networks [111] or random-forest-based doubly robust procedures [89], yet all of these methods require the time series to be strictly stationary, i.e. the joint probability distribution of the process is time-invariant, and assume a stochastic data-generating process. A complementary, model-free approach for single-molecule trajectories has been proposed by Berezhkovskii and Makarov [8]: by comparing transition-path probabilities $P(a\to b|x)$ with the value $1/4$ expected for one-dimensional Markov diffusion, it detects non-Markovian behavior without fitting the data to any model. However, this criterion is restricted to one-dimensional reaction coordinates and relies on equilibrium sampling. In the quantum-information context, Wolf et al. [103] introduced computable criteria and a continuous measure of Markovianity for quantum channels, deciding whether a given completely positive trace-preserving map is consistent with a Lindblad master equation. While this provides a rigorous framework for open quantum systems, it requires full quantum process tomography and operates within the Lindblad formalism; it is therefore not applicable to the reduced dynamics of a closed system. The trace-distance-based information-backflow criterion [15] was recently applied by Banerjee [5] to subsystem dynamics in a quenched Ising chain, detecting non-Markovianity of small subsystem reduced density matrices. However, this approach also relies on the open-system framework and scales exponentially with subsystem size, making it impractical beyond a few lattice sites. In contrast, neural ODEs impose no stationarity, probabilistic, or open-system assumptions. Moreover, because they learn a time-local vector field by construction, their predictive accuracy provides a continuous, physically interpretable measure of the degree of Markovianity, directly informing the development of closure functionals for reduced-order many-body methods.
@@
In this paper, we embark on a study of quantum many-body systems out of equilibrium, viewed through 2RDMs [34]. For this purpose, we investigate the dynamics of the paradigmatic Fermi-Hubbard model [80]. Depending on the parameters, it shows a broad range of dynamical regimes, reaching from weakly to strongly correlated, and from weak to strong excitation by tuning only two parameters. In the present work, we intentionally focus on small system sizes and comparatively simple lattice settings, since exact-reference data generation becomes rapidly more expensive with increasing system size and dimensionality. Within the computationally feasible regime, however, the study is already broad: we analyze over 300 distinct parameter configurations, each showing different dynamics and dynamical regimes. We demonstrate that neural ODEs offer interesting insights: We find indications that time-local reconstructions of the 3RDM may only be possible in certain regimes in the parameter space. The infusion of physics knowledge is neither straightforward nor does prior physical information necessarily lead to better results.
@@
The structure of the paper is as follows: In Sec. 2 we briefly introduce the quenched Fermi-Hubbard model, which will be the system under investigation. We describe the equations of motion for the 2RDM and concisely revisit the cumulant reconstruction needed to close them. The neural ODE model used in this investigation is presented together with the data preprocessing. In Sec. 3, we first identify some structure in the parameter space of the Fermi-Hubbard model. This is used to investigate the performance of the neural ODE across the parameter plane. To improve the stability of the prediction, we enforce physical constraints in the loss. Last, we discuss the results and show perspectives for future investigations in Sec. 4.
@@ 2.1 Non-equilibrium Fermi-Hubbard model
We consider a one‑dimensional Fermi–Hubbard chain comprising $M_{s}$ physical sites together with two auxiliary sites labeled 0 and $M_{s}+1$ that are fixed to be unoccupied throughout the dynamics. Dirichlet boundary conditions are imposed at these auxiliary sites, a choice that mirrors the confinement used in quantum‑simulator experiments with trapped ultracold atoms (see, e.g., Ref. [9]). While Dirichlet boundaries are natural for the experimental setting, alternative boundary conditions (e.g., periodic) can be incorporated straightforwardly within the same formalism. The initial state is prepared in a harmonic trapping potential, which is abruptly switched off at $t=0$, after which the system evolves under the standard Fermi-Hubbard Hamiltonian [34]. The full Hamiltonian in second quantization is given by
@@
Here, $J$ denotes the hopping amplitude, $\langle i,j\rangle$ indicates nearest-neighbor pairs, $a_{i\sigma}^{(\dagger)}$ are the single-particle annihilation (creation) operators, $n_{i}^{\uparrow(\downarrow)}=a_{i\uparrow(\downarrow)}^{\dagger}a_{i\uparrow(\downarrow)}$ are the corresponding spin-resolved occupation number operators at site $i$, and $U$ characterizes the on-site interaction strength governing the interaction energy. The system before and after the quench can be seen in Fig. 1.
@@
For the initial trap, we use a harmonic potential
@@
which initially traps the system in the ground state of the harmonic trap and is turned off at $t=0$. This potential quench, parametrized by $V$, induces the dynamics. Although this is a quite simple model, it is still complex enough to show rich dynamics.
@@ 2.2 Time-evolution of the 2RDM
A $p$-particle reduced density matrix (pRDM) is obtained from the NN‑body state $|\Psi(t)\rangle$ by tracing over all but $p$ particles
$D_{1\ldots p}(t)=\binom{N}{p}p!\text{Tr}_{p+1\ldots N}|\Psi(t)\rangle\langle\Psi(t)|,$ (3)
with $\binom{N}{p}p!$ being the chosen normalization for the pRDMs. This particular normalization has the advantage that the cumulant decompositions (see Eq. 6 and Eq. 7 below) do not have any prefactors. Correspondingly, the 2RDM is given by
$D_{12}(t)=N(N-1)\text{Tr}_{3\ldots N}|\Psi(t)\rangle\langle\Psi(t)|\,.$ (4)
The second equation in the BBGKY hierarchy, i.e., the equation of motion for the 2RDM,
$i\partial_{t}D_{12}=[h_{1}+h_{2}+W_{12},D_{12}]+\text{Tr}_{3}[W_{13}+W_{23},D_{123}]\,$, (5)
contains the single‑particle Hamiltonian $h_{i}$, the two‑body interaction $W_{ij}$, and the 3RDM $D_{123}$. $\text{Tr}_{3}$ denotes the trace over the third particle. Closing Eq. (5) therefore requires a reconstruction functional $D^{\text{R}}_{123}[D_{12}]$. This leads to the TD2RDM theory introduced in Refs. [55, 56, 34].
In a total‑spin‑singlet configuration, as considered here, the evaluation of Eq. (5) requires only the propagation of the mixed‑spin block $\langle j_{1}\uparrow j_{2}\downarrow|D_{12}(t)|i_{1}\uparrow i_{2}\downarrow\rangle=D^{j_{1}\uparrow j_{2}\downarrow}_{i_{1}\uparrow i_{2}\downarrow}(t)$, with $|i_{1}\uparrow i_{2}\downarrow\rangle$ two-particle product states of spin orbitals, since it already encodes the complete information of the full 2RDM $D_{12}(t)$ [55]. For completeness, we note that elements of the 2RDM can also be obtained in the second quantization formalism through $D^{j_{1}\uparrow j_{2}\downarrow}_{i_{1}\uparrow i_{2}\downarrow}(t)=\langle\Psi(t)|a_{i_{1}\uparrow}^{\dagger}a_{i_{2}\downarrow}^{\dagger}a_{j_{2}\downarrow}a_{j_{1}\uparrow}|\Psi(t)\rangle$. All other spin components can be recovered from this block by straightforward exchange or spin‑flip symmetry relations.